# LW-FedSSL: Resource-efficient Layer-wise Federated Self-supervised Learning

### v4

@@ 4.1 Experimental setup
4.2 Performance and Resource Comparison
Table 2 presents a comparison of the resource requirements and performance across different training approaches. [^2] We report mean per-class accuracy for the Caltech-101 dataset. For resource requirements, we compare the computational and communication resources needed for a single FL client. Additionally, we demonstrate that our proposed methods, LW-FedSSL and Prog-FedSSL, can be seamlessly integrated with various SSL techniques. As shown in Table 2, we integrate MoCoV3 [11], BYOL [16], and SimCLR [1] with our methods. For clarity, we group similar methods based on the SSL technique used. Fig. 5 plots the resource consumption for a client, illustrating computational resources (i.e., memory and FLOPs) across local epochs and communication cost (download + upload) across FL rounds. For FLOPs calculation, we consider only a single input sample.
The results in Table 2 show that LW-FedSSL significantly reduces resource requirements. Specifically, when using MoCoV3, LW-FedSSL achieves a 3.34×3.34$\times$3.34 × reduction in memory usage, a 4.20×4.20$\times$4.20 × fewer computational operations (GFLOPs), and a 5.07×5.07$\times$5.07 × reduction in communication costs compared to conventional FedMoCoV3. Additionally, Table 2 demonstrates that LW-FedSSL also maintains comparable performance to its end-to-end training counterparts. While Prog-FedSSL is not as resource-efficient as LW-FedSSL, it still consumes fewer GFLOPs and communication resources compared to FedMoCoV3. Furthermore, Prog-FedSSL achieves better performance in most cases, suggesting that progressive training can also benefit self-supervised learning, aligning with the findings for supervised learning in [24]. Fig. 5 illustrates that the resource consumption for LW-FedSSL remains relatively flat across local epochs and FL rounds due to the focus on training a single layer at each stage. Meanwhile, the resource consumption for Prog-FedSSL gradually increases as the number of trainable layers grows with each stage, eventually matching that of conventional FedMoCoV3.
Comm. (GB)
FedMoCoV3 8.40 (1.00×)
LW-FedSSL (ours) 3.35 (0.40×)
Prog-FedSSL (ours) 5.04 (0.60×)
Table 6: Resource requirements and performance comparison when integrating Depth Dropout (DD), Proximal Alignment (PA), and Representation Alignment (RA) into LW-FedSSL and Prog-FedSSL.
@@ 4.2 Performance and Resource Comparison
LW-FedSSL + DD [25] 2.49 (0.95×) 111 (0.80×) 1.66 (1.00×) 84.10 60.78 40.18 48.90 58.49
LW-FedSSL + PA [35] 2.63 (1.01×) 139 (1.00×) 1.66 (1.00×) 84.46 61.74 40.72 44.31 57.81
LW-FedSSL + RA [36] 2.86 (1.10×) 318 (2.29×) 1.66 (1.00×) 84.31 61.69 40.18 45.29 57.87
@@
Prog-FedSSL + DD [25] 5.24 (0.60×) 233 (0.73×) 3.97 (0.79×) 86.69 64.41 41.59 63.46 64.04
Prog-FedSSL + PA [35] 8.72 (1.00×) 318 (1.00×) 5.04 (1.00×) 84.71 62.09 38.65 56.35 60.45
Prog-FedSSL + RA [36] 9.00 (1.04×) 496 (1.56×) 5.04 (1.00×) 87.05 66.05 42.37 62.59 64.52
4.3 Worst-case Communication Scenario for LW-FedSSL
By default, we assume that clients only need to download the layer $L_{s}$ at each FL round for LW-FedSSL. This assumption holds only if all clients participate in every FL round without dropping out or no new clients join. Otherwise, new clients or those that previously dropped out would need to download the latest model $F_{s}$ (i.e., $[L_{1},\dots,L_{s}]$) instead of just $L_{s}$. Therefore, we compare the worst-case communication cost for LW-FedSSL in Table 5, where a client downloads $F_{s}$ at every FL round. Note that client still needs to upload only $L_{s}$ after local training. The results show that even in the worst-case scenario, LW-FedSSL significantly reduces communication costs compared to end-to-end FedMoCoV3.
4.4 Number of Layers Added per Stage
4.5 ResNet-18
To demonstrate that our proposed approaches can also work well with different model architectures, we use ResNet-18 [37] as the encoder $F$ in Table 4. We set the number of stages to $S=4$, with each stage sequentially adding a new ResNet block to the encoder. The total number of FL rounds is set to $R=120$, with each stage $s$ allocated $R_{s}=30$ rounds. Whenever a new ResNet block is added and the output dimension of the encoder changes, we add a temporary linear layer to reshape the output to 512. This reshaping layer is discarded at the end of each stage. The results in Table 4 demonstrate that our approaches also perform well with ResNet-18.
4.6 Integration of Additional Mechanisms
In both LW-FedSSL and Prog-FedSSL, each stage can be viewed as an independent FL process. At the beginning of each stage, some layers retain weights from the previous stage, but otherwise, the FL process within each stage remains independent. This modularity allows easy integration of existing FL mechanisms into LW-FedSSL and Prog-FedSSL. To demonstrate this flexibility, we incorporate depth dropout (DD) [25], a technique that selectively drops frozen layers during training; alignment using the proximal term (PA) [35], which constrains local updates to prevent drastic weight divergence across clients; and a similar alignment mechanism using representation (RA) [36]. Table 6 presents the evaluation results, including resource requirements and downstream performance across multiple datasets.
@@ 4.6 Integration of Additional Mechanisms
(a) PathMNIST
(b) BloodMNIST
(c) DermaMNIST
(d) OrganAMNIST
4.7 Partial Client Participation
@@
FedMoCoV3 33.56 27.24 27.85 26.97
LW-FedSSL (ours) 40.03 30.29 30.78 30.31
Prog-FedSSL (ours) 27.39 24.63 24.24 25.42
@@
4.8 Medical Data
We evaluate our proposed approaches on MedMNIST [38], a collection of diverse biomedical image datasets. Specifically, we use four subsets within MedMNIST: PathMNIST, BloodMNIST, DermaMNIST, and OrganAMNIST. Sample images from each dataset are shown in Fig. 7. Each dataset contains three splits: train, validation, and test. We use the train split to construct the local datasets for FL clients, the validation split as the downstream dataset for fine-tuning, and the test split for final evaluation. To introduce data heterogeneity, we partition the data among clients using the Dirichlet distribution [39] with a concentration parameter of $\beta=5$. We use an image size of $64\times 64$ and a batch size of 128 for all experiments. In Table 10, we present resource requirements only for BloodMNIST, as the experimental settings—such as batch size, input dimensions, and model architecture—are consistent across all datasets, resulting in similar resource requirements. In Table 9, we evaluate the performance of our proposed methods on the medical datasets. For all datasets, we report mean per-class accuracy.
Table 10: Resource requirements for the BloodMNIST dataset.
Resource Requirements (BloodMNIST)
Memory
GFLOPs
Comm.
FedMoCoV3 11.48 (1.00×) 2705 (1.00×) 8.45 (1.00×)
LW-FedSSL (ours) 3.10 (0.27×) 640 (0.24×) 1.66 (0.20×)
Prog-FedSSL (ours) 11.46 (1.00×) 1468 (0.54×) 5.09 (0.60×)
@@ 4.8 Medical Data
Memory (GB) GFLOPs Comm. (GB) Precision Recall F1 Accuracy
Image Size: $64\times 64$, Patch Size: $4\times 4$, Number of Patches: 256
4.9 Batch Size
We evaluate the impact of batch size on memory consumption using the BloodMNIST dataset. We consider batch sizes of 64, 128, 256, and 512 while all other settings remain the same as described in Section 4.8. Fig. 8 shows the peak memory usage across different batch sizes, while Table 12 presents the evaluation results. We only report memory usage in Fig. 8, as the FLOPs and communication costs across different batch sizes remain similar to those presented in Table 10.333This is because we consider only a single input sample for FLOPs calculations, and communication cost is independent of the training batch size. Given that Prog-FedSSL does not involve frozen layers, its peak memory requirement—which occurs at the final stage $S$—is equivalent to that of end-to-end FedMoCoV3. Peak memory requirements for both FedMoCoV3 and Prog-FedSSL sharply rise as the batch size grows, while those of LW-FedSSL remain relatively flat. Many SSL approaches that rely on contrastive loss require large batch sizes to leverage a greater number of negative samples [1]. The low-memory footprint of LW-FedSSL can accommodate large-batch training, making it more scalable in resource-constrained environments.
@@ 4.10 Impact of Input Dimensions
4.11 Data Heterogeneity
Data heterogeneity is one of the main challenges in an FL environment. The concentration parameter $\beta$ in the Dirichlet distribution can be used to determine the strength of data heterogeneity, with a lower $\beta$ value indicating a higher degree of heterogeneity. In Table 13, we conduct experiments using different levels of data heterogeneity by setting $\beta$ values to 50, 5, and 0.5.  [^4] The resulting client data distributions are visualized in Fig. 10. All other experimental settings are kept the same as in Section 4.8. Table 13 shows that all approaches are relatively robust to different $\beta$ values. This observation aligns with prior studies [8] that highlight the robustness of SSL-based approaches to data heterogeneity in FL environments. This also indicates that LW-FedSSL and Prog-FedSSL exhibit robustness to data heterogeneity inherent in end-to-end SSL training.
@@ 5 Conclusion
Edge devices in distributed environments often struggle to meet the computational and communication demands of federated self-supervised learning. In response to these challenges, we propose a layer-wise training approach named LW-FedSSL, which allows FL clients to perform self-supervised learning with significantly improved resource efficiency. By dividing the training process into multiple stages and focusing on a subset of layers at each stage, LW-FedSSL effectively reduces resource consumption. When applied to MoCoV3, LW-FedSSL requires 3.34×3.34$\times$3.34 × less memory, consumes 4.20×4.20$\times$4.20 × fewer GFLOPs, and reduces total transmission costs (download + upload) by 5.07×5.07$\times$5.07 ×, while maintaining comparable performance to its end-to-end counterpart, FedMoCoV3. Furthermore, we explore a progressive training approach named Prog-FedSSL. Although Prog-FedSSL is less resource-efficient than LW-FedSSL, it remains more efficient than end-to-end training and achieves superior performance in most cases. Through extensive experiments across various FL settings, datasets, and ablations, we demonstrate the effectiveness of LW-FedSSL and Prog-FedSSL. The results highlight the potential of layer-wise and progressive training strategies for enhancing the scalability and practicality of self-supervised learning in resource-constrained FL environments.
Appendix A Additional Details
A.1 Details on FLOPs Calculation
@@
Here, we discuss the details on FLOPs calculation used in our experiments. In layer-wise training, calculating FLOPs for inactive (frozen) layers involves simply considering FLOPs associated with the forward pass. However, for active layers, the calculation must account for both the forward and backward passes. While numerous works have addressed FLOPs calculation for the forward pass (inference), limited studies are available on the practice of FLOPs calculation for the backward pass. Notably, existing related works [24, 25] also do not provide specific details on how FLOPs are computed. Some studies [40, 41, 42, 43] suggest that the number of operations in a backward pass of a neural network is often twice that of a forward pass. Following these studies, we adopt a backward-forward ratio of 2:1 to calculate the total FLOPs for active layers. We use the FlopCountAnalysis tool in the fvcore [44] library for calculating the FLOPs of the forward pass. We only consider a single input sample for FLOPs calculation.
A.2 Data Distribution
@@
12:         $\mathcal{L}\leftarrow\ell_{\text{con}}(q_{1},k_{2},\tau)+\ell_{\text{con}}(q_{2},k_{1},\tau)$
@@
Figure 9: Prog-FedSSL: Local training process across different stages for the $n$-th client. At the beginning of each stage $s\in[1,S]$, a new layer $L_{s}$ is sequentially added to the encoder $F_{(s-1)}$, increasing its depth. During stage $s$, Prog-FedSSL updates all existing layers (i.e., $[L_{1},\dots,L_{s}]$) within $F_{s}$.
(a) $\beta=50$
(b) $\beta=5$
(c) $\beta=0.5$

### v5

@@ 4.3 Worst-case Communication Scenario for LW-FedSSL
4.4 Number of Layers Added per Stage
4.5 ResNet-18
To demonstrate that our proposed approaches can also work well with different model architectures, we use ResNet-18 [he2016deep] as the encoder $F$ in Table 5. We set the number of stages to $S=4$, with each stage sequentially adding a new ResNet block to the encoder. The total number of FL rounds is set to $R=120$, with each stage $s$ allocated $R_{s}=30$ rounds. Whenever a new ResNet block is added and the output dimension of the encoder changes, we add a temporary linear layer to reshape the output to 512. This reshaping layer is discarded at the end of each stage. The results in Table 5 demonstrate that our approaches also perform well with ResNet-18.
@@ 4.5 ResNet-18
4.6 Integration of Additional Mechanisms
In both LW-FedSSL and Prog-FedSSL, each stage can be viewed as an independent FL process. At the beginning of each stage, some layers retain weights from the previous stage, but apart from that, the FL process within each stage remains independent. This modularity allows easy integration of existing FL mechanisms into LW-FedSSL and Prog-FedSSL. To demonstrate this flexibility, we incorporate depth dropout (DD) [pengfei2023towards], a technique that selectively drops frozen layers during training; alignment using the proximal term (PA) [li2020federated], which constrains local updates to prevent drastic weight divergence across clients; and a similar alignment mechanism using representation (RA) [Li_2021_CVPR]. Table 6 presents the evaluation results, including resource requirements and downstream performance across multiple datasets.
Table 6: Resource requirements and performance comparison when integrating Depth Dropout ($D$), Proximal Alignment (PA), and Representation Alignment (RA) into LW-FedSSL and Prog-FedSSL.
@@ 4.6 Integration of Additional Mechanisms
LW-FedSSL + DD [pengfei2023towards] 2.49 (0.95×) 111 (0.80×) 1.66 (1.00×) 84.10 60.78 40.18 48.90 58.49
LW-FedSSL + PA [li2020federated] 2.63 (1.01×) 139 (1.00×) 1.66 (1.00×) 84.46 61.74 40.72 44.31 57.81
LW-FedSSL + RA [Li_2021_CVPR] 2.86 (1.10×) 318 (2.29×) 1.66 (1.00×) 84.31 61.69 40.18 45.29 57.87
@@
Prog-FedSSL + DD [pengfei2023towards] 5.24 (0.60×) 233 (0.73×) 3.97 (0.79×) 86.69 64.41 41.59 63.46 64.04
Prog-FedSSL + PA [li2020federated] 8.72 (1.00×) 318 (1.00×) 5.04 (1.00×) 84.71 62.09 38.65 56.35 60.45
Prog-FedSSL + RA [Li_2021_CVPR] 9.00 (1.04×) 496 (1.56×) 5.04 (1.00×) 87.05 66.05 42.37 62.59 64.52
4.7 Partial Client Participation
@@
4.8 Medical Data
We evaluate our proposed approaches on MedMNIST [medmnistv2], a collection of diverse biomedical image datasets. Specifically, we use four subsets within MedMNIST: PathMNIST, BloodMNIST, DermaMNIST, and OrganAMNIST. Sample images from each dataset are shown in Fig. 7. Each dataset contains three splits: train, validation, and test. We use the train split to construct the local datasets for FL clients, the validation split as the downstream dataset for fine-tuning, and the test split for final evaluation. To introduce data heterogeneity, we partition the data among clients using the Dirichlet distribution [ferguson1973bayesian] with a concentration parameter of $\beta=5$. We use an image size of $64\times 64$ and a batch size of 128 for all experiments. For DermaMNIST, we use a 3-layer classifier head with a hidden dimension of 256 and fine-tune for 100 epochs. All other settings follow the default configuration. In Table 9, we present resource requirements only for BloodMNIST, as the experimental settings—such as batch size, input dimensions, and model architecture—are consistent across all datasets, resulting in similar resource requirements. In Table 10, we evaluate the performance of our proposed methods on the medical datasets. For all datasets, we report mean per-class accuracy.
Table 9: Resource requirements for the BloodMNIST dataset.
Resource Requirements (BloodMNIST)
Memory
GFLOPs
Comm.
FedMoCoV3 11.48 (1.00×) 2705 (1.00×) 8.45 (1.00×)
LW-FedSSL (ours) 3.10 (0.27×) 640 (0.24×) 1.66 (0.20×)
Prog-FedSSL (ours) 11.46 (1.00×) 1468 (0.54×) 5.09 (0.60×)
@@ 4.8 Medical Data
FedMoCoV3 50.63 48.27 48.27 45.76
LW-FedSSL (ours) 51.33 47.55 48.15 46.06
Prog-FedSSL (ours) 51.73 49.38 49.15 47.72
@@
(a) PathMNIST
(b) BloodMNIST
(c) DermaMNIST
(d) OrganAMNIST
4.9 Impact of Batch Size Variation
We evaluate the impact of batch size on memory consumption using the BloodMNIST dataset. We consider batch sizes of 64, 128, 256, and 512 while all other settings remain the same as described in Section 4.8. Fig. 8 shows the peak memory usage across different batch sizes, while Table 11 presents the evaluation results. We only report memory usage in Fig. 8, as the FLOPs and communication costs across different batch sizes remain similar to those presented in Table 9.333This is because we consider only a single input sample for FLOPs calculations, and communication cost is independent of the training batch size. Given that Prog-FedSSL does not involve frozen layers, its peak memory requirement—which occurs at the final stage SS—is equivalent to that of end-to-end FedMoCoV3. Peak memory requirements for both FedMoCoV3 and Prog-FedSSL sharply rise as the batch size grows, while those of LW-FedSSL remain relatively flat. Many SSL approaches that rely on contrastive loss require large batch sizes to leverage a greater number of negative samples [chen2020simple]. The low-memory footprint of LW-FedSSL can accommodate large-batch training, making it more scalable in resource-constrained environments.
@@ 4.10 Impact of Input Dimensions
4.11 Data Heterogeneity
Data heterogeneity is one of the main challenges in an FL environment. The concentration parameter $\beta$ in the Dirichlet distribution can be used to determine the strength of data heterogeneity, with a lower $\beta$ value indicating a higher degree of heterogeneity. In Table 13, we conduct experiments using different levels of data heterogeneity by setting $\beta$ values to 50, 5, and 0.5.  [^4] The resulting client data distributions are visualized in Fig. 10. All other experimental settings are kept the same as in Section 4.8. Table 13 shows that all approaches are relatively robust to different $\beta$ values. This observation aligns with prior studies [wang2023does] that highlight the robustness of SSL-based approaches to data heterogeneity in FL environments. This also indicates that LW-FedSSL and Prog-FedSSL exhibit robustness to data heterogeneity inherent in end-to-end SSL training.
Memory (GB) GFLOPs Comm. (GB) Precision Recall F1 Accuracy
Image Size: $64\times 64$, Patch Size: $4\times 4$, Number of Patches: 256
@@ 5 Conclusion
Edge devices in distributed environments often struggle to meet the computational and communication demands of federated self-supervised learning. In response to these challenges, we propose a layer-wise training approach named LW-FedSSL, which allows FL clients to perform self-supervised learning with significantly improved resource efficiency. By dividing the training process into multiple stages and focusing on a subset of layers at each stage, LW-FedSSL effectively reduces resource consumption. When applied to MoCoV3, LW-FedSSL requires $3.34\times$ less memory, consumes $4.20\times$ fewer GFLOPs, and reduces total transmission costs (download + upload) by $5.07\times$, while maintaining comparable performance to its end-to-end counterpart, FedMoCoV3. Furthermore, we explore a progressive training approach named Prog-FedSSL. Although Prog-FedSSL is less resource-efficient than LW-FedSSL, it remains more efficient than end-to-end training and achieves superior performance in most cases. Through extensive experiments across various FL settings, datasets, and ablations, we demonstrate the effectiveness of LW-FedSSL and Prog-FedSSL. The results highlight the potential of layer-wise and progressive training strategies for enhancing the scalability and practicality of self-supervised learning in resource-constrained FL environments.
Appendix A Convergence Behavior
To examine the convergence behavior of our proposed strategies, we plot the average client loss at each communication round for both LW-FedSSL and Prog-FedSSL. Fig. 9 shows the results for two settings: $S=2$ and $S=12$ (the default). For LW-FedSSL, we observe that both settings converge over the course of training. However, with $S=12$, the loss remains slightly higher than that of FedMoCoV3, while reducing the number of stages to $S=2$ results in convergence behavior closer to FedMoCoV3. This is expected since FedMoCoV3 can be regarded as training with one stage (i.e., $S=1$). For Prog-FedSSL, both stage settings demonstrate convergence patterns similar to FedMoCoV3.
(a) LW-FedSSL.
(b) Prog-FedSSL.
Figure 9: Average client loss over communication rounds for (a) LW-FedSSL and (b) Prog-FedSSL, using stage settings $S=2$ and $S=12$ (default), compared with end-to-end FedMoCoV3. LW-FedSSL converges in both settings, with $S=2$ exhibiting convergence behavior closer to FedMoCoV3. Prog-FedSSL converges similarly to FedMoCoV3 in both settings.
Appendix B Additional Details
B.1 Details on FLOPs Calculation
@@
Here, we discuss the details on FLOPs calculation used in our experiments. In layer-wise training, calculating FLOPs for inactive (frozen) layers involves simply considering FLOPs associated with the forward pass. However, for active layers, the calculation must account for both the forward and backward passes. While numerous works have addressed FLOPs calculation for the forward pass (inference), limited studies are available on the practice of FLOPs calculation for the backward pass. Notably, existing related works [wang2022progfed, pengfei2023towards] also do not provide specific details on how FLOPs are computed. Some studies [estimate, whattheback, aiandcompute, epoch2021backwardforwardFLOPratio] suggest that the number of operations in a backward pass of a neural network is often twice that of a forward pass. Following these studies, we adopt a backward-forward ratio of 2:1 to calculate the total FLOPs for active layers. We use the FlopCountAnalysis tool in the fvcore [fvcore] library for calculating the FLOPs of the forward pass. We only consider a single input sample for FLOPs calculation.
(a) $\beta=50$
(b) $\beta=5$
(c) $\beta=0.5$
Figure 11: Prog-FedSSL: Local training process across different stages for the $n$-th client. At the beginning of each stage $s\in[1,S]$, a new layer $L_{s}$ is sequentially added to the encoder $F_{(s-1)}$, increasing its depth. During stage $s$, Prog-FedSSL updates all existing layers (i.e., $L_{1},\dots,L_{s}]$) within $F_{s}$.
B.2 Data Distribution
@@
12:   $\mathcal{L}\leftarrow\ell_{\text{con}}(q_{1},k_{2},\tau)+\ell_{\text{con}}(q_{2},k_{1},\tau)$ $\triangleright$ Local loss